OpenStudy (anonymous):

s(s+4)=0 solve for s

OpenStudy (anonymous):

its s(s+4)=96

OpenStudy (anonymous):

s=0 or x=-4

OpenStudy (anonymous):

s=0 or s=-4

OpenStudy (anonymous):

s^2+4s=96 s^2+4s-96=0 now solve for x

OpenStudy (anonymous):

sorry solve for s

OpenStudy (anonymous):

96? from where you grapped that?

OpenStudy (anonymous):

that was the problem i typed it wrong

OpenStudy (anonymous):

it's already factored equation, which means you solve directly.

OpenStudy (anonymous):

ok, np.